OpenStudy (anonymous):

integrate by partial fractions, not sure,. i want to integrate 1/((x^2*(x^2-15)) should i do partial fraction 1/((x^2*(x^2-15)) = A / x + B/x^2 + C/(x-sqrt 15) + D/(x+sqrt 15) , or 1/((x^2*(x^2-15)) = A / x + B/x^2 + Cx+d / ( x^2 - 15)

OpenStudy (amistre64):

oy... if you hit the enter key it spreads this out to where it is more readable.... just a thought

OpenStudy (anonymous):

the x^2 - 15 is reducible , but im not sure if youre supposed to use that theorem

OpenStudy (anonymous):

enter key?

OpenStudy (amistre64):

the key on the keyboard that makes this happen ;)

OpenStudy (anonymous):

you want me to use the enter key more often?

OpenStudy (amistre64):

yes, difference of squares that part and you should have 3 sets to figure out

OpenStudy (anonymous):

but you get radicals

OpenStudy (amistre64):

radicals are fine.... right?

OpenStudy (anonymous):

are they ? how come the other way works just as well?

OpenStudy (amistre64):

which other way can you decompose the fraction?

OpenStudy (anonymous):

1/((x^2*(x^2-15)) = A / x + B/x^2 + (Cx+d) / ( x^2 - 15)

OpenStudy (anonymous):

technically the quadratic is reducible

OpenStudy (amistre64):

it could be a fluke if it does work out....

OpenStudy (amistre64):

or it could be a possibility for solutions that i am unaware of as yet :)

OpenStudy (anonymous):

no it isnt a fluke, in general one need not reduce a quadratic

OpenStudy (amistre64):

in general you should, to account for multiples i think

OpenStudy (amistre64):

if it aint got multiples, then you should be fine i spose

OpenStudy (amistre64):

right?

OpenStudy (anonymous):

yes it usually makes it easier to reduce the quadratic, except in the case where there are radical linear factors

OpenStudy (anonymous):

it can have multiples, theres no problem as far as i can see

OpenStudy (amistre64):

without doing a ton of them to generalize it, id have to throw it in the hmmmmm box lol

OpenStudy (anonymous):

say integral ( 1 / (x(x+1)^2) = A/x + (Bx+C) / (x+1)^2

OpenStudy (anonymous):

its a little weird, might be easier if i wrote 1/( x ( x^2+2x+1)) = A/x + (Bx+c) / (x^2 + 2x + 1)

OpenStudy (anonymous):

sometimes it is difficult to factor a quadratic, right?

OpenStudy (anonymous):

ok , i should get into more accurate terms. sometimes it is reducible over reals, not rationals

OpenStudy (amistre64):

factoring a quadratic is fairly easy; might get complicated solutions, but farily easy to do

OpenStudy (anonymous):

yes but doing partial fractions with irrational roots is tough

OpenStudy (amistre64):

and i have heard that you can use the complex solutions as well with no ill effects on the decomp

OpenStudy (anonymous):

ok , so youre so confident, try to do parfractions on 1 / ( x^2(x^2-15) , give it a go

OpenStudy (anonymous):

i was stuck

OpenStudy (anonymous):

1/((x^2*(x^2-15)) = A / x + B/x^2 + C/(x-sqrt 15) + D/(x+sqrt 15)

OpenStudy (amistre64):

confident? lol ; just stubborn

OpenStudy (anonymous):

it can even get stranger, this is an easy irrational solution problem, you can have say A/ ( x - sqrt 3/2 ) for instance

OpenStudy (anonymous):

this is why calculus in textbooks is generally contrived , or idealized problems. in the real world you have these irrational roots, etc

OpenStudy (anonymous):

theorem time, one sec

OpenStudy (amistre64):

1 a b c d ----------- = --- + ----- + --------- + --------- x^2(x^2-15) x x^2 x-sqrt(15) x+sqrt(15) is the 'proper' set up right?

OpenStudy (anonymous):

yes looks good

OpenStudy (anonymous):

Personally, I'd leave the last fraction's denominator without doing difference of two squares.

OpenStudy (anonymous):

1 / [(x-a)^n (x^2+bx+c)^n] = A1/(x-a) + A2/ ( x-a)^2 + ... An/(x-a)^n + (B1x+C1)/ (x^2+bx+c) + (B2x + C2) / (x^2+bx+c)^2 + ... + (Bn x + Cn ) / (x^2+bx+c)^n

OpenStudy (anonymous):

so the question is, do the quadratics have to be irreducible, i dont think so

OpenStudy (amistre64):

1 a b c d ----------- = --- + ----- + --------- + --------- x^2(x^2-15) x x^2 x-sqrt(15) x+sqrt(15) a(x^2)(x-sqrt(15))(x+sqrt(15))) b(x)(x-sqrt(15))(x+sqrt(15))) c(x)( x^2)(x+sqrt(15)) d(x)(x^2)(x-sqrt(15))

OpenStudy (anonymous):

MOST calculus books and websites say the quadratic is irreducible though

OpenStudy (anonymous):

irreducible over the reals i imagine. wikipedia is better, they say irreducible over the field

OpenStudy (anonymous):

whatever field you have in mine

OpenStudy (anonymous):

For this particular problem: \[\frac{1}{y(y-15)} \equiv \frac{1}{15(y-15)} - \frac{1}{15y}\] \[\text{Let } y = x^2\] ...

OpenStudy (anonymous):

newton, be careful though

OpenStudy (anonymous):

if you are substituting a variable, you will need to do 2x dx = dy

OpenStudy (anonymous):

so you might want to just change it back. thats an identity

OpenStudy (anonymous):

i should write a book on calculus pitfalls, since i keep finding them :)

OpenStudy (anonymous):

newtons, so just change y back to x^2

OpenStudy (anonymous):

Sorry, in case it wasn't clear, it leads to: \[\frac{1}{x^2(x^2-15)} \equiv \frac{1}{15(x^2-15)} + \frac{1}{15x^2} \] I would integrate this.

OpenStudy (anonymous):

otherwise you have 2x dx = dy , dx = dy / ( 2x ) = dy / ( 2 sqrt y )

OpenStudy (anonymous):

correct, i thought you were doing a change of variables

OpenStudy (amistre64):

1 = a(x^2)(x-sqrt(15))(x+sqrt(15))) + b(x)(x-sqrt(15))(x+sqrt(15))) + c(x)( x^2)(x+sqrt(15)) +d(x)(x^2)(x-sqrt(15)) when x = sqrt(15) we get 1 = c(sqrt(15))(15)(2sqrt(15)) 1 = c(15)(2)(15) 1 = c(225)(2) = c450 c = 1/450

OpenStudy (anonymous):

amistre, you get medal for persistence :)

OpenStudy (anonymous):

but its much easier if we dont split the quadratic

OpenStudy (anonymous):

I was only doing that to show that you don't need a 1/x fraction and 1/x^2 , as x^2 is the only term. I wasn't integrating with a different variable.

OpenStudy (amistre64):

yay!! lol :) can I stop now lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait that doesnt work though

OpenStudy (anonymous):

I mixed up a minus and plus sign ¬_¬

OpenStudy (anonymous):

its incorrect to do 1 / ( x^2 (x^2-15 ) = A / x^2 + b / (x^2-15) , thats incorrect

OpenStudy (anonymous):

you wont be able to find A and B

OpenStudy (anonymous):

I just did?

OpenStudy (anonymous):

you used a simpler case though

OpenStudy (anonymous):

Multiply it out, plz.

OpenStudy (anonymous):

1 = A(x^2 - 15) + B x^2

OpenStudy (anonymous):

, 1 = Ax^2 + bx^2 - 15

OpenStudy (anonymous):

(the + the last fraction should be minus, as in the first time, typo)

OpenStudy (anonymous):

1 = Ax^2 + Bx62 - 15 A

OpenStudy (anonymous):

hmmmm,

OpenStudy (anonymous):

but the book says it should be A/ x + B/x^2 + C x+d / (x^2 - 15)

OpenStudy (anonymous):

well this approach works as well

OpenStudy (anonymous):

You can do it if you want, but you'll find A = 0, so it doesn't really matter.

OpenStudy (anonymous):

right, so come the theorem doesnt work in these websites

OpenStudy (anonymous):

as far as repeated roots go

OpenStudy (anonymous):

for instance you didnt use CX+D, i think its because of that y = x^2 business,

OpenStudy (anonymous):

There are only terms in x^2 in this case, it wouldn't always work. The general method is probably more ... general.

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so my question, my original question, hmmmm, do the quadratics have to be irreducible in the general case

OpenStudy (anonymous):

ok for instance, 1 / (( x-2)^2 ( x^2-15))

OpenStudy (anonymous):

1 / (( x-2)^2 ( x^2-15)) = A / (x-2) + b / (x-2)^2 + (cx+d) / (x^2 - 15 ) ?

OpenStudy (anonymous):

I'd say it's probably easier to leave it as the irreducible (without using difference of two squares) quadratic with Cx + d. I guess you could split up this factor, but it's unnecessary (as far as I know)

OpenStudy (anonymous):

Yes, that.

OpenStudy (anonymous):

but its not technically irreducible

OpenStudy (anonymous):

i can even go so far as to do this

OpenStudy (anonymous):

1 / (( x-2)^2 ( x^2-15)) = (Ax+b) / (x-2)^2 + (cx+d) / (x^2 - 15 )

OpenStudy (anonymous):

no chat room

OpenStudy (anonymous):

sorry, not for you

OpenStudy (anonymous):

so this whole irreducible business is unnecessary

OpenStudy (anonymous):

1 = a*x^3-15*a*x+b*x^2-15*b+c*x^3-4*c*x^2+4*c*x+d*x^2-4*d*x+4*d

OpenStudy (anonymous):

ok im going to solve an easier one

OpenStudy (anonymous):

int 1/ ((x-2)(x^2-15))

OpenStudy (anonymous):

1/ ((x-2)(x^2-15)) = A/ (x-2) + (Bx+c) / ( x^ 2 - 15)